# When Do Credal Sets Stabilize? Fixed-Point Theorems for Credal Set Updates

### v1

@@ Appendix B Additional Result: Finitely Additive Case
In this section, we generalize Corollary 2.2 to the finitely additive case; as we shall see, we need to pay a price for this extension.
@@
A direct consequence of Lemma 7 is the following, which is a version of Corollary 2.2 in the finitely additive case.
@@
By Lemma 7, we know that $\mathcal{X}$ finite implies that $(\mathscr{C}^{\text{fa}},d_{H})$ is a complete metric space. The result, then, follows from the Boyd-Wong fixed point theorem (Boyd and Wong,, 1969, Theorem 1). ∎
As we can see, in the finitely additive setting, it is not enough for $\mathcal{X}$ to be Polish, as pointed out in Remark 2. To accommodate for finite continuity, we must require that $\mathcal{X}$ is finite. This means that we move from a topological to a set-theoretic structure requirement for $\mathcal{X}$. The latter condition is of course more restrictive than the ones we had in Section 3.

### v2

@@ Appendix D Additional Result: Finitely Additive Case
In this section, we generalize Corollary 3 to the finitely additive case; as we shall see, we need to pay a price for this extension.
@@
A direct consequence of Lemma 11 is the following, which is a version of Corollary 3 in the finitely additive case.
@@
By Lemma 11, we know that $\mathcal{X}$ finite implies that $(\mathscr{C}^{\text{fa}},d_{H})$ is a complete metric space. The result, then, follows from Boyd-Wong’s fixed point theorem, Theorem 8 in Appendix A.5. ∎
As we can see, in the finitely additive setting, it is not enough for $\mathcal{X}$ to be Polish, as in Theorem 3. To accommodate for finite continuity, we must require that $\mathcal{X}$ is finite. This means that we move from a topological to a set-theoretic structure requirement for $\mathcal{X}$. The latter condition is of course more restrictive than the ones we had in Section 3.